OpenStudy (anonymous):

Find the given term of each arithmetic sequence. 1/4,3/4,5/4,7/4,...;15th term

OpenStudy (amistre64):

since it says its arithmetic; that makes life easier. A{n} = A{n-1} + d = A{n-2} + d + d = A{n-3} + d + d + d = A{n-3} + 3d = A{n-r} + rd when r = n-1, we can get to A1 A{n} = A{n-(n-1)} + (n-1)d = A1 + (n-1)d this is the general form of an arithmetic sequance

OpenStudy (amistre64):

in this case A1 = 1/4 and the "d"ifference between each term is 2/4 An = 1/4 + 2/4 (n-1)

OpenStudy (anonymous):

so i am still confused